OpenStudy (tay1lynn):

Solve for x: 3 over 4 x + 5 over 8 = 4x @agent0smith

OpenStudy (amit1235813):

Step 1: Multiply both sides by 8

OpenStudy (skullpatrol):

Step 0: translate the words into symbols

OpenStudy (tyrenhere):

^^

OpenStudy (amit1235813):

lol yup I assumed Step 0

OpenStudy (skullpatrol):

Math is about learning how to walk on your own: step by step ;-)

OpenStudy (tyrenhere):

@tay1lynn Do you want the answer?

OpenStudy (tay1lynn):

yes please

OpenStudy (tay1lynn):

@TyrenHere

OpenStudy (tyrenhere):

@quickstudent He's good with stuff like this.

OpenStudy (skullpatrol):

Change: 3 over 4 x + 5 over 8 = 4x Into symbols please :)

OpenStudy (quickstudent):

Me? haha, I'm a girl. OK, let me see how I can help:)

OpenStudy (tyrenhere):

Oh! Im sorry.

OpenStudy (tay1lynn):

thank you

OpenStudy (quickstudent):

(3/4)x + (5/8) = 4x So here's the equation.

OpenStudy (skullpatrol):

@tay1lynn do you want to learn?

OpenStudy (quickstudent):

Here's the first step. You need to isolate the x. (3/4)x + (5/8) = 4x (3/4)x + (5/8) - (3/4)x = 4x - (3/4)x (5/8) = 3 1/4 Do you understand this?

OpenStudy (tay1lynn):

yes thank you

OpenStudy (quickstudent):

By the way, if I make any mistake anywhere, anybody is welcomed to correct me:)

OpenStudy (quickstudent):

woops, typo, missed the x. Here it is (5/8) = 3 1/4x

OpenStudy (quickstudent):

Or maybe even simpler: (5/8) = (3 1/4)x

OpenStudy (quickstudent):

Now do you know what to do next? (remember, we are isolating the x)

OpenStudy (quickstudent):

As we did before, we are working to isolate the x, so here is how you would go further: (5/8)/(3 1/4) = (3 1/4)x/(3 1/4) (5/26) = x Do you get it? If you have any trouble with the steps, please ask:)